# Reducing Optimism Bias in Incomplete Cooperative Games

### v1

@@ 1. Introduction
In practice, acquiring even a single value for a single coalition can be a resource-intensive endeavor. Take, for instance, the realm of machine learning, where determining the value of a feature subset that represents a coalition in the celebrated explainable approach SHAP Lundberg and Lee (2017) corresponds to retraining an entire model, consuming time and computational resources. What is more, these contributions can have ripple effects on subsequent financial outlays, such as acquiring new training samples. In the corporate world, estimating an employee’s contribution to collective performance may facilitate their fair evaluation Murphy and Cleveland (1995), but obtaining the value of a coalition may involve the intricate process of rearranging teams of employees, incurring operational costs and potentially causing disruptions.
@@ 2. Preliminaries
$\phi_{i}(v)\coloneqq\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!(|N|-|S|-1)!}{|N|!}\bigl{(}v(S\cup\{i\})-v(S)\bigr{)},$ (4)
The Shapley value is efficient, which captures the fact that exactly $v(N)$ is distributed among the players. Specifically,
@@
Another important property of the Shapley value is linearity,
$\phi(\alpha v+\beta w)=\alpha\cdot\phi(v)+\beta\cdot\phi(w).$ (6)
The characteristic function $v$ is represented by $2^{n}$ real values. In many applications, however, obtaining all the values might be too expensive. To deal with this problem, $\mathcal{K}\subseteq 2^{N}$ is introduced in the definition of an incomplete game which represents only coalitions with known value. In our application, we model obtaining new information about the unknown (but well defined) values by extending the set $\mathcal{K}$. Therefore, we view $\mathcal{K}$ as a ‘masking set’ applied to some complete game and use the characteristic function of that complete game in our definition [^2] This is in contrast with the standard definition where only $v:\mathcal{K}\to\mathbb{R}$ is used.. Another distinction from the standard definition is that we require knowledge of at least minimal information333 In the standard definition, only $\emptyset\in\mathcal{K}$ is required Masuya and Inuiguchi (2016); Bok and Černý (2023).
@@
If further properties of the underlying game are assumed, one may impose restrictions on values of $S\notin\mathcal{K}$ even though they are not known exactly. This is quite natural in many situations and common throughout the literature to assume superadditivity of the underlying game Masuya and Inuiguchi (2016); Peleg and Sudhölter (2007). Under this assumption and based on partial knowledge represented by $\mathcal{K}$, one can describe the set of candidates for the underlying game. The set consists of extensions of the partial set function, which satisfy superadditivity.
@@ 3. Optimism Bias and Its Minimization
From (9), (10), it is immediate $(N,v_{i})$ extends $(N,\mathcal{K},v)$. Further, Theorem 3 in Masuya and Inuiguchi (2016) states that any cooperative game with a characteristic function defined for a non-empty coalition $S$ as
is superadditive. We obtain the result by setting $S=\{i\}$. ∎
@@
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{i\in N}\phi_{i}(v_{i})-v(N)=\sum_{S\subseteq N}\alpha_{S}\Delta_{\hat{\mathcal{K}}}(S)-v(N),$ (16)
@@
which is exactly $\phi_{i}(v_{i})$. The proof of the second equality is merely a technicality left for the Appendix H. ∎
@@ 3.1. Principal’s Optimization Problems
To minimize the utopian gap, we hence assume the existence of a non-affiliated party who we refer to as the principal. Her task is to choose which coalitions should be investigated to reduce the utopian gap the most. We further assume the principal holds a level of expertise that guides the selection process. This expertise is formalized as a prior distribution over a set of potential characteristic functions. For example, in a medical context, a doctor acting as the principal might seek to assess a patient’s response to a combination of drugs (represented as a coalition) and base their expectations on past clinical experience. Similarly, a machine learning engineer could rely on their prior knowledge of feature importance gained from previous problem-solving experiences. Consequently, we assume that each problem instance can be viewed as a sample drawn from a known prior distribution, denoted as $\mathcal{F}$. To put bluntly, the principal is aware of the prior distribution, but not of the specific instance of the underlying game.
@@
where $\mathcal{K}_{\tau}$, $\tau\leq t$ is such that $\mathcal{K}_{\tau}=\mathcal{K}_{\tau-1}\cup\{S_{\tau}\}$, $S_{\tau}=\pi(N,\mathcal{K}_{\tau-1},v)$ where $\pi$ is a policy function, which chooses $S_{\tau}$ based on the known values of $v$, i.e. values of coalitions in $\mathcal{K}_{\tau-1}$.
@@ 3.2.1. Offline Algorithms
¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus\mathcal{K}_{0}$ $G\leftarrow\left\{\right\}$   // trajectories and their average gap
@@
2       $\mu\leftarrow 0$ for $j\in\{1,\dots,\kappa\}$ do // approximate expectation
3             $v\sim\mathcal{F}$ $\mu\leftarrow\mu+\mathcal{G}_{(N,v)}(\mathcal{S})$
4      $\mu\leftarrow\mu/\kappa$ $G\left[\mathcal{S}\right]\leftarrow\mu$
$\left\{S_{i}\right\}_{i=1}^{t}\leftarrow\operatornamewithlimits{argmin}_{\mathcal{S}\subseteq\overline{}\mathcal{K}:\left\lvert\mathcal{S}\right\rvert=t}G\left[\mathcal{S}\right]$ return $\{S_{i}\}_{i=1}^{t}$
@@ 4      $\mu\leftarrow\mu/\kappa$ $G\left[\mathcal{S}\right]\leftarrow\mu$
$\{S_{i}\}_{i=1}^{t-1}\leftarrow$ Offline Greedy($\mathcal{F},t-1$) ¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus(\mathcal{K}_{0}\cup\left\{S_{i}\right\}_{i=1}^{t-1})$ $G\leftarrow\left\{\right\}$   // trajectories & their average gap
@@
2       $\mu\leftarrow 0$ for $j\in\{1,\dots,\kappa\}$ do // approximate expectation
3             $v\sim\mathcal{F}$ $\mu\leftarrow\mu+\mathcal{G}_{(N,v)}(\left\{S_{i}\right\}_{i=1}^{t-1}\cup\left\{S\right\})$
4      $\mu\leftarrow\mu/\kappa$ $G\left[S\right]\leftarrow\mu$
$S_{t}\leftarrow\operatornamewithlimits{argmin}_{S\in\overline{}\mathcal{K}}G\left[S\right]$ return $\{S_{i}\}_{i=1}^{t}$
@@ 3.2.2. Online Algorithm
As previously mentioned, the greedy algorithm is significantly more computationally efficient, with a linear complexity in the number of coalitions, while the optimal variant exhibits an exponential time complexity. A question which naturally arises is under which conditions the local greedy search can reliably yield a globally optimal solution. While our empirical observations indicate similar performance between both approaches, we arrive at negative conclusions regarding this inquiry. To illustrate, we provide a simple example wherein the local search method fails to identify the global optimum. In fact, in this particular scenario, the distribution $\mathcal{F}$ is extremely simplistic, comprising only a single element [^4] Hence, in this scenario, both offline and online approaches yield identical results..
[…]
Let ($(\{o,w_{1},w_{2},w_{3}\},v)$ be the underlying cooperative game with a characteristic function $v$ defined as
[…]
Starting with the minimal information, the greedy strategy proceeds to select coalitions $\left\{w_{1},w_{2},w_{3}\right\}$, $\left\{o,w_{1},w_{2}\right\}$, and $\left\{o,w_{3}\right\}$, in the initial three steps. However, at fourth step, the greedy strategy selects $\left\{o,w_{1},w_{2}\right\}$, while the trajectory that minimizes the gap entails choosing all coalitions of size three. Consequently, the greedy strategy achieves a larger utopian gap. It is worth noting that the greedy strategy strictly prefers its choice at step $t=3$. See Appendix C for more details.
It is known that, if the optimized function is supermodular, the locally optimal steps are guaranteed to yield a $(1-1/e)$-approximation of the global optimum (Nemhauser et al., 1978, Proposition 3.4). In our case, the utopian gap is supermodular for all games with $|N|\leq 4$.
@@
However, when $|N|\geq 5$, the gap is not supermodular already for the game in Example 3.2.2 (even for $|N|=5$). Additionally, analogous results can be derived for games in other significant subsets of superadditive games, such as totally monotonic games (a subset of convex games with relevance in measure theory Grabisch (2016)), symmetric games (often studied for their robust properties), or subgraph games (games with a compact structure represented by $O(n^{2})$ values Chalkiadakis et al. (2012)). The proof of Proposition 3.2.2 and the counter-examples are available in Appendix G. Despite these findings, empirically, the greedy algorithm demonstrates similar performance to its optimal counterpart, as illustrated in the following section.
@@ 4. Empirical Evaluation
Finally, we demonstrate the performance of our algorithms on practical examples. We outline the domains used for our evaluation, introduce baseline methods for comparison, detail the algorithmic setups, and conduct a comprehensive analysis of the gathered results. All the details that could not be accommodated within the main text have been diligently addressed in the relevant appendices.
@@ 4.1. Experimental Domains
In words, each worker contributes equally to the value of a coalition as long as he has somewhere to work. It is easy to verify that the resulting game is superadditive. We denote the distribution of factory games of $n$ players where the owner is selected uniformly at random as factory$(n)$.
@@
We conduct further studies on several other classes of cooperative games as well, and present these results in Appendix D. We observe similar trends as those presented here in the main text.
@@ 4.3. Experimental Setup
During the training phase, we execute the PPO algorithm for a total of 2,000,000 time-steps, which, in the case of 12-step instances, translates to 40 iterations. In each iteration, we repeatedly collect 4,096 trajectories by sampling batches from the distribution $\mathcal{F}$. Following each iteration, we optimize the PPO surrogate objective for 10 epochs. To ensure uniformity, we normalize the input values $v(S)$ to a unit interval, as elaborated in Appendix F. For the two optimal algorithms, we employ $\kappa=10$ samples to estimate the mean value in Eq. (21).
See Appendix B for more details.
Among the algorithms we have proposed, both Offline Optimal and Oracle Optimal come with significantly higher computational costs [^5] Their complexity scales as $\mathcal{O}(2^{2^{n}})$ in terms of the number of players. when compared to their more computationally efficient greedy counterparts. To provide some perspective, our evaluation for $n=5$ with twelve steps and ten samples necessitated approximately 60 hours to finish. Extrapolating, we estimate that completing the evaluations for the remaining steps in this setup, across all ten samples, would require roughly 200 hours. Furthermore, extending these experiments to $n=6$ would demand over 100 years.
@@
All experiments are conducted on a computational cluster with AMD EPYC 7532 CPUs running at 2.4 GHz. When running algorithms on five players games, we utilize 15 cores and 12 GB of RAM. The code was implemented in Python 3.10 using pytorch 2.0, stable_baselines3 2.0, and gymnasium 0.28. Additional details are presented in Appendix B. The code is available in the supplementary material and will be released upon acceptance.
@@ 5. Conclusion
We posit that our approach is not confined solely to the boundaries of cooperative game theory and can be extended to more general set functions. This extension assumes the existence of criteria involving functions [^6] In this work, it is the Shapley value. that map the powerset of a set $N$ to subspaces of dimensions linear in the size of $N$. Still in the context of games, we see applications of our approach within the SHAP Lundberg and Lee (2017) framework, as the original formulation can only handle comparatively small models. Finally, we would also like to strengthen our result by providing guarantees by working within the regret minimization framework.
Acknowledgments
The authors would like to thank Martin Loebl and Milan Hladík for their insightful comments. David Sychrovský and Martin Černý received support from the Charles University Grant Agency (GAUK 206523). This work was supported by the Czech Science Foundation grant no. GA22-26655S, the CoSP Project grant no. 823748, the Charles University project UNCE 24/SCI/008, and the Grant Agency. Computational resources were supplied by the project e-Infrastruktura CZ (e-INFRA LM2018140) provided within the program Projects of Large Research, Development and Innovations Infrastructures.

### v3

@@ 1. Introduction
In practice, acquiring even a single value for a single coalition can be a resource-intensive endeavor. Take, for instance, the realm of machine learning, where determining the value of a feature subset that represents a coalition in the celebrated explainable approach SHAP Lundberg and Lee (2017) corresponds to retraining an entire model, consuming time and computational resources. What is more, these contributions can have ripple effects on subsequent financial outlays, such as acquiring new training samples.
In the corporate world, estimating an employee’s contribution to collective performance may facilitate their fair evaluation Murphy and Cleveland (1995), but obtaining the value of a coalition may involve the intricate process of rearranging teams of employees, incurring operational costs and potentially causing disruptions.
@@ 2. Preliminaries
$\phi_{i}(v)\coloneqq\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!(|N|-|S|-1)!}{|N|!}\bigl{(}v(S\cup\{i\})-v(S)\bigr{)}.$ (4)
The Shapley value is linear and efficient, which captures the fact that exactly $v(N)$ is distributed among the players. Specifically,
@@
The characteristic function $v$ is represented by $2^{n}$ real values. In many applications, however, obtaining all the values might be too expensive. To deal with this problem, $\mathcal{K}\subseteq 2^{N}$ is introduced in the definition of an incomplete game which represents only coalitions with known value. In our application, we model obtaining new information about the unknown (but well defined) values by extending the set $\mathcal{K}$. Therefore, we view $\mathcal{K}$ as a ‘masking set’ applied to some complete game and use the characteristic function of that complete game in our definition. [^2] This is in contrast with the standard definition where only $v:\mathcal{K}\to\mathbb{R}$ is used. Another distinction from the standard definition is that we require knowledge of at least minimal information333 In the standard definition, only $\emptyset\in\mathcal{K}$ is required Masuya and Inuiguchi (2016); Bok and Černý (2023).
@@
If further properties of the underlying game are assumed, one may impose restrictions on values of $S\notin\mathcal{K}$ even though they are not known exactly. The property we consider is superadditivity, which is natural in many situations and common throughout the literature on exchange economies, cost-distribution problems, or supermodular optimization Masuya and Inuiguchi (2016); Peleg and Sudhölter (2007). [^4] The immediacy of this property is not always guaranteed. For instance, SHAP may not consistently yield a superadditive game. However, within this particular context, the property is observed, for example, in uncorrelated models within an ensemble. Under this assumption and based on partial knowledge represented by $\mathcal{K}$, one can describe the set of candidates for the underlying game. The set consists of extensions of the partial set function, which satisfy superadditivity.
@@ 3. Optimism Bias and Its Minimization
From (8), (9), it is immediate $(N,v_{i})$ extends $(N,\mathcal{K},v)$. Further, as mentioned earlier, Theorem 3 in Masuya and Inuiguchi (2016) states that any cooperative game with a characteristic function (11) is superadditive. We obtain the result by setting $S=\{i\}$. ∎
@@
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{i\in N}\phi_{i}(v_{i})-v(N)=\sum_{S\subseteq N}\alpha_{S}\Delta_{\hat{\mathcal{K}}}(S),$ (14)
@@
which is exactly $\phi_{i}(v_{i})$. The proof of the second equality is merely a technicality left for the Appendix G. ∎
@@ 3.1. Principal’s Optimization Problems
To minimize the utopian gap, we hence assume the existence of a non-affiliated party who we refer to as the principal. Her task is to choose which coalitions should be investigated to reduce the utopian gap the most. We further assume the principal holds a level of expertise that guides the selection process. This expertise is formalized as a prior distribution over a set of potential characteristic functions. For example, in a medical context, a doctor acting as the principal might seek to assess a patient’s response to a combination of drugs (represented as a coalition) and base their expectations on past clinical experience. Similarly, a machine learning engineer could rely on their prior knowledge of feature importance gained from previous problem-solving experiences. Consequently, we assume that each problem instance can be viewed as a sample drawn from this known prior distribution, denoted as $\mathcal{F}$. To put bluntly, the principal is aware of the prior distribution, but not of the specific instance of the underlying game.
@@
where $\mathcal{K}_{\tau}$, $\tau\leq t$ is such that $\mathcal{K}_{\tau}=\mathcal{K}_{\tau-1}\cup\{S_{\tau}\}$, $S_{\tau}=\pi(N,\mathcal{K}_{\tau-1},v)$, and $\pi$ is a policy function that chooses $S_{\tau}$ based on the known values of $v$, i.e. values of coalitions in $\mathcal{K}_{\tau-1}$.
@@ 3.2.1. Offline Algorithms
1 ¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus\mathcal{K}_{0}$
$G\leftarrow\left\{\right\}$
  // trajectories and their average gap
@@
3       $\mu\leftarrow 0$
4       for $j\in\{1,\dots,\kappa\}$ do // approximate expectation
5             $v\sim\mathcal{F}$
6             $\mu\leftarrow\mu+\mathcal{G}_{(N,v)}(\mathcal{S})$
7
8      $\mu\leftarrow\mu/\kappa$
9       $G\left[\mathcal{S}\right]\leftarrow\mu$
10$\left\{S_{i}\right\}_{i=1}^{t}\leftarrow\operatornamewithlimits{argmin}_{\mathcal{S}\subseteq\overline{}\mathcal{K}:\left\lvert\mathcal{S}\right\rvert=t}G\left[\mathcal{S}\right]$
11 return $\{S_{i}\}_{i=1}^{t}$
@@ 9       $G\left[\mathcal{S}\right]\leftarrow\mu$
1 $\{S_{i}\}_{i=1}^{t-1}\leftarrow$ Offline Greedy($\mathcal{F},t-1$)
2 ¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus(\mathcal{K}_{0}\cup\left\{S_{i}\right\}_{i=1}^{t-1})$
$G\leftarrow\left\{\right\}$
  // trajectories & their average gap
@@ 1 $\{S_{i}\}_{i=1}^{t-1}\leftarrow$ Offline Greedy($\mathcal{F},t-1$)
4       $\mu\leftarrow 0$
5       for $j\in\{1,\dots,\kappa\}$ do // approximate expectation
6             $v\sim\mathcal{F}$
7             $\mu\leftarrow\mu+\mathcal{G}_{(N,v)}(\left\{S_{i}\right\}_{i=1}^{t-1}\cup\left\{S\right\})$
8
9      $\mu\leftarrow\mu/\kappa$
10       $G\left[S\right]\leftarrow\mu$
[…]
12 return $\{S_{i}\}_{i=1}^{t}$
@@ 3.2.2. Online Algorithm
As previously mentioned, the greedy algorithm is significantly more computationally efficient, with a linear complexity in the number of coalitions, while the optimal variant exhibits an exponential time complexity. A question which naturally arises is under which conditions the local greedy search can reliably yield a globally optimal solution. While our empirical observations indicate similar performance between both approaches, the greedy method is not optimal in general. To illustrate, we provide a simple example in Appendix C, where the local search fails to identify the optimum. However, it is known that, if the optimized function is supermodular, the locally optimal steps are guaranteed to yield a $(1-1/e)$-approximation of the global optimum (George et al., 1978, Proposition 3.4). In our case, the utopian gap is supermodular for all small games.
@@
Proof Sketch.
or $\lvert N\rvert=3$, the gap is modular due to invariant bounds post-revelation. For $|N|=4$, the supermodularity is demonstrated through a more technical case analysis. The full proof can be found in Appendix H.1. ∎
However, already for $|N|=5$, it can be shown the gap is not supermodular for several significant subsets of superadditive games, such as totally monotonic games (a subset of convex games, which under the name belief measures forms the foundation on the evidence theory Grabisch (2016); Shafer (1976)), symmetric games (often studied for their robust properties with respect to solution concepts Peleg and Sudhölter (2007)), or graph games (games with a compact structure represented by $O(n^{2})$ values with application in scenarios where only bilateral collaboration between players occurs Chalkiadakis et al. (2012)).
Observation.
For $|N|=5$, there exist a totally monotonic game, convex game, symmetric superadditive game, and graph game for which the utopian gap is not supermodular.
For even larger games, we can formualate a precise criterion to ascertain the non-supermodularity of the utopian gap.
[…]
For $|N|\geq 6$, the gap of an incomplete $\mathbb{S}^{n}$-extendable game is not supermodular if there exist $i,j,k,l\in N$ such that $\tilde{v}(ij)\leq\tilde{v}(jk)\leq\tilde{v}(kl)$, where $\tilde{v}(S)=v(S)-\sum_{i\in S}v(\{i\})$, and
$\tilde{v}(kl)<\left[{\binom{n}{2}}{\binom{n}{\lfloor n/2\rfloor}}^{-1}\left(2^{n-3}-n+2\right)-1\right]\tilde{v}(ij).$ (20)
Proof Sketch.
We rephrase the supermodularity condition (3) in relation to the utopian gap, and demonstrate that this condition is never met when Eq. (20) holds for the pairs in some $\{i,j,k,l\}$. The full proof can be found in Appendix H.3. ∎
This criterion allows to construct examples also for all earlier mentioned classes of games. The proofs and examples are available in Appendix H. Despite these findings, empirically, the greedy algorithm demonstrates similar performance to its optimal counterpart, as illustrated in the following section.
@@ 4. Empirical Evaluation
Finally, we demonstrate the performance of our algorithms on practical examples. We outline the domains used for the evaluation, introduce baseline methods for comparison, detail the algorithmic setups, and conduct a comprehensive analysis of the gathered results. All the details that could not be accommodated within the main text have been addressed in the relevant appendices.
@@ 4.1. Experimental Domains
In words, each worker contributes equally to the value of a coalition as long as he has somewhere to work. It is easy to verify that the resulting game is superadditive. We denote the distribution of factory games of $n$ players where the owner is selected uniformly at random as factory$(n)$. We show that the gap is not supermodular already for factory($5$) in Appendix C.
@@
We conduct further experiments on several other classes of cooperative games as well, and present these results in Appendix D. We observe similar trends as those presented here in the main text.
@@ 4.3. Experimental Setup
During the training phase, we execute the PPO algorithm for a total of 2,000,000 time-steps, which, in the case of 12-step instances, translates to 40 iterations. In each iteration, we repeatedly collect 4,096 trajectories by sampling batches from the distribution $\mathcal{F}$. Following each iteration, we optimize the PPO surrogate objective for 10 epochs. To ensure uniformity, we normalize the input values $v(S)$ to a unit interval, as elaborated in Appendix F. For the two optimal algorithms, we employ $\kappa=10$ samples to estimate the mean value in Eq. (19). See Appendix B for more details.
Among the algorithms we have proposed, both Offline Optimal and Oracle Optimal come with significantly higher computational costs (their complexity scales as $\mathcal{O}(2^{2^{n}})$) when compared to their more computationally efficient greedy counterparts (which scale $\mathcal{O}({2^{n}})$). To provide some perspective, our evaluation for $n=5$ with twelve steps and ten samples necessitated approximately 60 hours to finish. Extrapolating, we estimate that completing the evaluations for the remaining steps in this setup, across all ten samples, would require roughly 200 hours. Furthermore, extending these experiments to $n=6$ would demand over 100 years.
@@
All experiments are conducted on a computational cluster with AMD EPYC 7532 CPUs running at 2.4 GHz. When running algorithms on five players games, we utilize 15 cores and 12 GB of RAM. The code was implemented in Python 3.10 using pytorch 2.0, stable_baselines3 2.0, and gymnasium 0.28. Additional details are presented in Appendix B, and is available at GitHub Úradník (2024).
@@ 5. Conclusion
We posit that our approach is not confined solely to the boundaries of cooperative game theory and can be extended to more general set functions. This extension assumes the existence of criteria involving functions (in this work, it is the Shapley value). that map the powerset of a set $N$ to subspaces of dimensions linear in the size of $N$. Still in the context of games, we see applications of our approach within the SHAP Lundberg and Lee (2017) framework, as the original formulation can only handle comparatively small models. Finally, we would also like to strengthen our result by providing guarantees by working within the regret minimization framework.
{acks}
The authors would like to thank Martin Loebl and Milan Hladík for their insightful comments. David Sychrovský and Martin Černý received support from the Charles University Grant Agency (GAUK 206523). This work was supported by the Czech Science Foundation grant no. GA22-26655S, the CoSP Project grant no. 823748, and the Charles University project UNCE 24/SCI/008. Computational resources were supplied by the project e-Infrastruktura CZ (e-INFRA LM2018140) provided within the program Projects of Large Research, Development and Innovations Infrastructures.
@@ Appendix C Example: Local Optimum is not Global Optimum
C.1. Non-Supermodularity for $|N|>4$
In order for the utopian gap to be supermodular, it needs to satisfy $\forall S,Z\in 2^{N}$, and $\forall\mathcal{K}\subseteq 2^{N},\mathcal{K}_{0}\subseteq\mathcal{K}$
However, when the owner is player 1 and one selects $\mathcal{K}=\mathcal{K}_{0}\cup\{\{1,2,3\},\{1,4\}\}$, $S=\{1,2\}$, and $Z=\{1,2,3,5\}$, the condition reads $7.3-7.9-8.1+8.6=-0.1<0$, which shows factory ($5$) is not supermodular. Similar argument can be made for $n>5$.